A solution contains 4.9 grams of sugar dissolved in 23.6 grams of water. What is the mass % of the solution?


1
Expert's answer
2021-04-05T07:46:43-0400

"\omega=\dfrac{m_{sugar}}{m_{solution}}=\dfrac{4.9 g}{4.9 g + 23.6 g}=0.1719=17.19\%"